Use the method of completing the square to find the standard form of the quadratic function.
f(x) = x^2 + 6x − 1
y=
x=
(x,y)=

Answer by lwsshak3(11628)

Use the method of completing the square to find the standard form of the quadratic function.
f(x) = x^2 + 6x − 1
= x^2 + 6x-1=0
x^2 +6x=1
x^2+6x+9=1+9
(x+3)^2=10
(x+3=±√10
x=-3±√10